OpenStudy (anonymous):

The snack bar at your school has added sushi to its menu. The ingredients for one roll include sushi rice, seaweed sheets, cucumbers, cream cheese, and 3oz of smoked salmon. One roll can be cut into 8 servings. Write an expression for the amount of salmon needed to make s serving of sushi. How much salmon is needed to make 16 servings? 24 servings? 80 servings? 100 servings?

OpenStudy (anonymous):

Here's the full thing! ^^

OpenStudy (anonymous):

thank you for clicking on this.

OpenStudy (anonymous):

1 roll makes 8 servings, and each roll contains 3 oz of salmon. So 3 oz makes 8 servings. 6 oz makes 16 servings. 9 oz makes 24 servings. 3/8 = x/80 (cross multiply) 8x = 240 x = 30 oz makes 80 servings 3/8 = x/100 (cross multiply) 8x = 300 x = 37.5 oz makes 100 servings. I am not sure what the expression would be. You can set them up like a proportion....3/8 = x oz/servings. Then cross multiply and solve for x (number of ounces.)

OpenStudy (anonymous):

@phi can you help here...please

OpenStudy (anonymous):

Noooo, I got it! Thank you for your help!!!

OpenStudy (anonymous):

I think you divide the number of servings by 8 then take that answer and multiply it by 3. I think

OpenStudy (anonymous):

but any ideas on how to make into a table?

OpenStudy (anonymous):

well....let me see... 3 oz 8 servings 6 16 9 24 12 32 15 40 18 48 21 56 24 64 does that make sense ?

OpenStudy (anonymous):

you divide the number of servings by 8, then multiply that by 3 to get the oz's

OpenStudy (anonymous):

it makes a little sense, but it says to get the amount of salmon needed to make the servings? not the rolls right?

OpenStudy (phi):

to do these type problems I write ratios along with the label for example, they say 3 oz per 1 roll \[ \frac{ 3 \text{ oz}}{1 \text { roll} }\] 1 roll per 8 servings \[ \frac{ 1 \text { roll} }{ 8 \text{ serv}}\] now if we want x amount of salmon (ounces) per s servings \[ \frac{ \text{ x oz}}{\text{ s serv}} \] I would match up my ratios so that when we multiply them we get oz per serv notice \[ \frac{ 3 \text{ oz}}{1 \text { roll} } \cdot \frac{ 1 \text { roll} }{ 8 \text{ serv}}\] does the trick. we cancel the roll \[ \frac{ 3 \text{ oz}}{\cancel{1 \text { roll}} } \cdot \frac{ \cancel{1 \text { roll}} }{ 8 \text{ serv}}\] we can say \[ \frac{ \text{ x oz}}{\text{ s serv}} = \frac{ 3 \text{ oz}}{8 \text{ serv}}\] now multipy both sides by x serve to get \[ \text{ x oz} = \frac{3\cdot s\text{ oz}}{8} \] notice the serv cancelled. that is your equation

OpenStudy (phi):

*now multipy both sides by s serve to get

OpenStudy (anonymous):

yes....the amount of salmon used can be found by taking the number of servings divided by 8, then multiply that by 3 . lets take 48 servings. Divide that by 8 and you get 6. Now multiply that by 3 and you get 18. That is the oz's of salmon used for 48 servings.

OpenStudy (anonymous):

phi gave you the equation.....that should help you bunches.

OpenStudy (anonymous):

Ahhhhhhhhhh, I get it!!! like,  do that for each one of them and I can answer them all! is that right?

OpenStudy (phi):

yes, but when you get to 100, you will get fractions

OpenStudy (phi):

you should write it \[ \frac{48\cdot 3}{8} = 6 \cdot 3 = 18 \ oz\]

OpenStudy (anonymous):

I see what your talking about on the 100 one, is 75/ 2? Just to be sure that I'm doing it right?

OpenStudy (phi):

yes that looks good. but 75/2 is 37.5 which how I would write it.

OpenStudy (anonymous):

awwwwwww, you two! thank you two soo much!!! the both of you are great people for helping others with this stuff!!! @phi and @kelliegirl33 that's all the help I need, thank you!!! Actaully, one more question, feel free to call me stupid but how would you write a table for this problem???

OpenStudy (anonymous):

it's okay, I think I can figure out one for myself!! Thanks again!!!

OpenStudy (phi):

list the number of servings in the first column and the number of ounces of salmon in the 2nd column

OpenStudy (anonymous):

thank you!